OpenStudy (anonymous):

Given the linear characteristic polynomial of a differential equation, find the order. (L+3)^3(L^2-2L+3)=0

OpenStudy (amistre64):

order refers to what, the highest derivative?

OpenStudy (anonymous):

@amistre64 yes, so 5

OpenStudy (amistre64):

5 looks agreeable to me as well

OpenStudy (anonymous):

For the linear characteristic polynomial \((r+3)^3 (r^2-2r+3) \)

OpenStudy (anonymous):

@amistre64 How do I find the general solution from it?

OpenStudy (amistre64):

solve for L L+3 = 0 L^2 -2L + 3= 0 depending on what the L represents ... its either going to be a string of\[e^{Lx}~or~x^{L}\]

OpenStudy (anonymous):

@amistre64 L stands for lambda.

OpenStudy (amistre64):

\[(L+3)^3(L^2-2L+3)\] L^3+3L^2+3L+1 L^2-2L+3 ----------------- \(L^5+3L^4+3L^3+L^2\) \(-2L^4-6L^3-6L^2-2L\) \(3L^3+9L^2+9L+3\) ------------------------------ \(L^5+L^4+0L^3+4L^2+7L+3\) this could come from a setup: \[y^{(5)}+y^{(4)}+4y''+7y'+3y=0\]

OpenStudy (anonymous):

@amistre64 Thank you a lot, you're a lifesaver!

OpenStudy (amistre64):

the solution for this can come in a variety of forms: e^(Lx) is a common evaluation for y x^(L) is another common evaluation say y = x^m y' = m x^(m-1) y'' = m(m-1) x^(m-2) y''' = m(m-1)(m-2) x^(m-3) etc ... filling in the y''' parts and stripping out the x^{m-5} parts i believe leaves us with a similar characteristic equation

OpenStudy (amistre64):

lets assume e^(Lx) is our general solution format L=-3,-3,-3,1-i sqrt(2), 1+isqrt(2) when we have multiple roots such as those -3, we have to include some xs to ensure a linearly independant solution of all possible configuration: \[y=c_1e^{-3x}+c_2xe^{-3x}+c_3x^2e^{-3x}+c_4e^{1-i~\sqrt{2}}+c_5e^{1+i~\sqrt{2}}\] i believe those i parts can simplify to a cos sin setup